OpenStudy (solomonzelman):

@johnns12234, I can't show the latex and space needed in a msg, so I posted it like this.

OpenStudy (solomonzelman):

1.) Solve each system using substitution. Tell whether the system has one solution, infinitely many solutions, or no solution. -x + y = -13 3x - y =19

OpenStudy (anonymous):

Alright!

OpenStudy (solomonzelman):

-x + y = -13 3x - y =19 ------------ (add them) 2x + 0y = 6 2x=6 x=3 Solve for y -x + y = -13 -(3) + y = -13 -3+y=-13 y=-10 does it work? -(3) + (-10) = -13 ----> -3-10=-13 3(3) - (-10) =19 ----> 6-(-10)=19 ---> 6+10=19

OpenStudy (solomonzelman):

So yes, there is one solution, and it's extraneous (it works) x=3 y=-10 or (3,-10)

OpenStudy (anonymous):

Thank you so much! I have a couple more wannt help? haha

OpenStudy (anonymous):

I have you your medal!

OpenStudy (solomonzelman):

Oh, thank you for the medal. the main thing is that you get it, and I see you do...

OpenStudy (anonymous):

huh?

OpenStudy (solomonzelman):

you would probably nvm me on this one, whatever...

OpenStudy (anonymous):

Wait... im so confused

OpenStudy (solomonzelman):

I meant becuase Id get the "huh?" you would probably tell me "nvm"...

OpenStudy (anonymous):

okay....?

OpenStudy (solomonzelman):

Yeah, just forget out about it ....

OpenStudy (anonymous):

alright...